bran:

Write the equation of the line that contains the indicated pair of points. Express the final equation in standard form. (0, 0), (−2, 9)

Tranquility:

First find the slope of the two points If the two points are \( (x_1, y_1)\) and \( (x_2,y_2)\) then the slope is \(\Large \text{Slope} =m= \frac{y_2 - y_1 }{x_2 - x_1}\) Once you find the slope, you can put it in to this form \( y - y_1 = m(x-x_1)\) You can choose either point to plug in to \( (x_1, y_1)\) Once you get that, you just need to rearrange that equation to get it into standard form. Remember the standard form for a linear equation is Ax + By = C and A cannot be negative. Also remember that A, B, and C should all be integers so you might need to multiply it by some coefficient if you get fractions.